OpenStudy (anonymous):

what percent of 56 is 224

OpenStudy (anonymous):

224/56=x

OpenStudy (anonymous):

i mean switch that around 56/224= 0.25 answer 25%

OpenStudy (anonymous):

anything else? @dylanlaskey

OpenStudy (danjs):

What percent - x/100 of - multiply is - equals \[\frac{ x }{ 100 }*56 = 224\]

OpenStudy (anonymous):

what

OpenStudy (anonymous):

was that a question?

OpenStudy (danjs):

that is the answer, solve for x%

OpenStudy (anonymous):

or you can just do it my way

OpenStudy (danjs):

224 is 400% of 56

OpenStudy (anonymous):

like this: 56/224 not 224/56

OpenStudy (danjs):

no, you are doing the reverse - What percent of 224 is 56.

OpenStudy (danjs):

Follow the solution method i did to translate the sentence into an equation.

OpenStudy (anonymous):

maybe he typed it wrong?

OpenStudy (danjs):

"What percent of 56 is 224?" \[\frac{ x }{ 100 } * 56 = 224\] \[\frac{ x }{ 100 } = \frac{ 224 }{ 56 }\] 400%

OpenStudy (danjs):

To translate, What percent means x/100 Of means multiply is means equals